### v2

# On measures derived from orbital integrals

Abstract.
In this article we propose a novel framework to derive piecewise polynomial measures which result from invariant measures on adjoint orbits in the general context of compact and semi-simple Lie groups. The measures are computed from orbital integrals by a series of transformations on spaces of polynomials endowed with the apolar inner product.
Keywords. orbital integral, orbital measures, piecewise polynomial measure, apolar inner product, Fisher Bombieri inner product, reproducing kernel.
The author was supported by IAM-CONICET, grants PIP 2010-0757 (CONICET) and PICT 2010-2478 (ANPCyT)
1. Introduction
In this article we develop a novel approach to the derivation of some piecewise polynomial measures or Duistermat-Heckman type measures from Harish-Chandras (HC) orbital integral formula in the context of compact semi-simple Lie groups. These measures are derived from orbital measures, that is, invariant probability measures on adjoint orbits of compact semi-simple Lie groups. Specifically, we derive the pushforward of an orbital measure by the projection to the Cartan algebra, and we derive the radial part of the convolution of two orbital measures. These problems were previously addressed in [Z16, O13, F15] and in [F19, CMZ19, Z18]. They are probabilistic versions of the Schur-Horn theorem [H54] and Horn’s problem [H62]. Given hermitian matrices $X$ and $Y$ with fixed eigenvalues the Schur-Horn theorem characterizes the possible diagonal entries of $X$ as the convex closure of the permutation of eigenvalues, while Horn’s problem asks for a characterization of the possible eigenvalues of the sum $X+Y$. The possible eigenvalues are given by a set of linear inequalities and the problem was solved by Klyatchko in [K98].
The approach applies directly to compact semi-simple Lie groups and is based on operations which can be done in finite dimensional spaces of polynomials endowed with the apolar inner product, which is the Segal-Bargmann-Fock space inner product restricted to polynomials. This operation if performed to both sides of an equation where the Fourier-Laplace transform of the unknown measure is written as an exponential polynomial involving the discriminant. The framework in the article was motivated in part by [M25] where the Harish-Chandra-Itzykson-Zuber (HCIZ) integral formula was put in the context of Segal-Bargmann spaces, and also [PS09] where divided difference operators and their adjoints were used to solve algebraic problems. In this article we took finite degree parts of power series and performed operations on these polynomials. Proofs in which finite degree terms are considered separately and an exponential function is constructed at the end are not uncommon in related literature, see for example the character expansion in the HCIZ integral in [MS17] and the proof of the Duistermaat-Heckman localization formula [MS17].
The article is organized as follows. In Section 2 we present the necessary results to state and prove the main theorems of the article. In Section 3 we give a characterization of the pushforward of an orbital measure and in Section 4 we give a characterization of the radial part of the convolution of two orbital measures.
2. Preliminaries
We review the results which are necessary to prove the main results of the article, such as the Harish-Chandra integral formula and operators between spaces of polynomials endowed with the apolar inner product. Most results are known, to the best of our knowledge only Propositions 2.5, 2.7, 2.9 and 2.10 are new.
2.1. Orbital measures and the Harish-Chandra integral formula
Let $G$ be a compact, semisimple and connected Lie group with Lie algebra $\operatorname{\mathfrak{g}}$. The Killing form is the $\operatorname{Ad}$-invariant inner product in $\operatorname{\mathfrak{g}}$ and is denoted with $\langle\cdot,\cdot\rangle$, $\operatorname{\mathcal{W}}$ is the Weyl group, $\operatorname{\epsilon}(w)$ is the sign of $w\in\operatorname{\mathcal{W}}$, that is, $\operatorname{\epsilon}(w)=(-1)^{|w|}$ where $|w|$ is the number of reflections necessary to generate $w$. A compact Lie group admits a unique invariant probability measure called the Haar measure and the adjoint action induces a unique invariant probability measure $\nu_{a}$ on the orbits $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{G}(a)$ for $a\in\operatorname{\mathfrak{g}}$. In [HC57] Harish-Chandra proved a formula for orbital integrals in Lie algebras in the case of compact, connected and semi-simple Lie groups, see the expository article [M21] and the references therein. The orbital integral admits an expression as an exponential polynomial. It is usually written
$\Delta(x)\Delta(y)\int_{G}e^{\langle\operatorname{Ad}_{g}x,y\rangle}dg=\frac{\left[\Delta,\Delta\right]}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(x),y\rangle},$
where $x$ and $y$ are in a Cartan algebra $\operatorname{\mathfrak{t}}_{\operatorname{\mathbb{C}}}$ of $\operatorname{\mathfrak{g}}_{\operatorname{\mathbb{C}}}$. The discriminant
$\Delta(x)=\prod_{\alpha\in\Phi^{+}}\langle\alpha,x\rangle$
is the product of the positive roots $\Phi^{+}$ considered as linear functionals. The term $\left[\Delta,\Delta\right]$ is computed with the apolar inner product which we a are going to introduce in the next section. We usually consider the element $a\in\operatorname{\mathfrak{t}}$ of the orbit $\operatorname{\mathcal{O}}_{a}$ as fixed and we write when $\Delta(a)\neq 0$
(1) $\operatorname{\mathcal{F}}_{a}(x)=\int_{G}e^{\langle\operatorname{Ad}_{g}a,x\rangle}dg=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}$
as the Fourier-Laplace transform of the probability measure $\nu_{a}$ on the orbit $\operatorname{\mathcal{O}}_{a}$.
2.2. Spaces of polynomials and operators between them
Let $V$ be a vector space of dimension $n\in\operatorname{\mathbb{N}}$ with the inner product $\langle\cdot,\cdot\rangle$. This space has coordinates $(x_{1},\dots,x_{n})$ given by an orthonormal basis. We define the space of polynomials on $V$ with degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ as
$\operatorname{\mathcal{P}}^{k}(V)=\{f\in\operatorname{\mathcal{P}}(V):\deg(f)\leq k\}.$
This space is endowed with the apolar inner product
$\left[f,g\right]=f(\partial)g(x)|_{x=0}$
for $f,g\in\operatorname{\mathcal{P}}^{k}(V)$. Here $f(\partial)$ is the polynomial $f(\frac{\partial}{\partial x_{1}},\dots,\frac{\partial}{\partial x_{n}})$. This inner product is sometimes called Fisher or Bombieri inner product. If $(x_{1},\dots,x_{n})$ are coordinates given by an orthonormal basis of $V$ then the monomials given by $x^{\beta}/\sqrt{\beta!}$ are orthonormal, where $\beta\in\operatorname{\mathbb{N}}^{n}$ is a multi-index and $\beta!:=\beta_{1}!\beta_{2}!\dots\beta_{n}!$. Note that
$\left[x^{\beta},x^{\gamma}\right]=\begin{cases}\beta!&\text{if $\beta=\gamma$}\\ 0&\text{if $\beta\neq\gamma$}\end{cases}$
Let
$f(x)=\sum_{\beta}f_{\beta}x^{\beta}\qquad\mbox{and}\qquad g(x)=\sum_{\beta}g_{\beta}x^{\beta}$
be two polynomials on $V$ where $\beta$ runs on multi-indices, then
$\left[f,g\right]=f(\partial)g(x)|_{x=0}=\sum_{\beta}f_{\beta}g_{\beta}\beta!.$
From this formula it follows that the apolar inner product is symmetric.
Remark 2.1.
The apolar inner product is the Segal-Bargmann-Fock inner product [Ha00, N11] restricted to real polynomials. If $f,g$ are polynomials defined over the complex numbers then
$g^{*}(\partial)f(z)|_{z=0}=\frac{1}{\pi^{n}}\int_{\operatorname{\mathbb{C}}^{n}}f(z)\overline{g(z)}e^{-|z|^{2}}dz$
where $g^{*}(z)=\overline{g(\overline{z})}$.
A multivariate version of Taylor’s theorem applied to polynomials asserts that for a polynomial $f$ and $x,a\in V$
$f(x+a)=e^{\langle a,\partial\rangle}f(x)=\sum_{\beta}a^{\beta}\partial^{\beta}f(x).$
If for $a\in V$ we denote by $T_{a}f(x)=f(x+a)$ the translation operator on polynomials then $T_{a}=e^{\langle a,\partial\rangle}$ as operators on polynomials. In the space $\operatorname{\mathcal{P}}^{k}(V)$ we define for $a\in V$ the functions
$q^{k}_{a}(x)=\sum_{l=0}^{k}\langle x,a\rangle^{l}$
which are truncations of exponentials $e^{\langle x,a\rangle}$. We denote with $\operatorname{ev}_{a}(f)=f(a)$ the evaluation functionals at $a\in V$.
Proposition 2.2.
For $a\in V$ the functions $q_{a}^{k}$ has the reproducing property
$\left[f,q_{a}^{k}\right]=\operatorname{ev}_{a}(f)=f(a)$
for $f\in\operatorname{\mathcal{P}}^{k}(V).$
Proof.
For $a\in V$ and $f\in\operatorname{\mathcal{P}}^{k}(V)$
$\left[f,q_{a}^{k}\right]$ $=q_{a}^{k}(\partial)f(x)|_{x=0}=e^{\langle\partial,a\rangle}f(x)|_{x=0}$
$=f(x+a)|_{x=0}=f(a)$
where we used the definition of the the apolar inner product and Taylor’s theorem for polynomials. This property also follows from the reproducing property of Segal-Bargmann spaces restricted to polynomials, see Section $5$ in [Ha00]. ∎
Consider the case when the inner product space is endowed with a finite refection group $\operatorname{\mathcal{W}}$ which has an alternating character $\epsilon:\operatorname{\mathcal{W}}\to\{1,-1\}$. We can define the space of alternating polynomials
$\operatorname{\mathcal{P}}^{k}_{\operatorname{alt}}(V)=\{f\in\operatorname{\mathcal{P}}^{k}(V):f(w(x))=\epsilon(w)f(x)\mbox{ for all }w\in\operatorname{\mathcal{W}}\}$
and the space of symmetric polynomials
$\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(V)=\{f\in\operatorname{\mathcal{P}}^{k}(V):f(w(x))=f(x)\mbox{ for all }w\in\operatorname{\mathcal{W}}\}.$
In $\operatorname{\mathcal{P}}^{k}(V)$ we define the orthogonal projection onto the alternating polynomials $P_{\operatorname{alt}}$ and the orthogonal projection onto the symmetric $P_{\operatorname{sym}}$, they are given by
$P_{\operatorname{alt}}f(x)=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)f(w(x))\qquad\mbox{and}\qquad P_{\operatorname{sym}}f(x)=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}f(w(x)).$
Lemma 2.3.
For $f\in\operatorname{\mathcal{P}}^{k}(V)$ and $g\in\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(V)$ we have
$\left[P_{\operatorname{alt}}f,g\right]=\left[f,g\right]$
Proof.
This follows from $\left[P_{\operatorname{alt}}f,g\right]=\left[f,P_{\operatorname{alt}}g\right]=\left[f,g\right]$, since $P_{\operatorname{alt}}^{*}=P_{\operatorname{alt}}$ and projections fix the vectors in their range. ∎
We define for $a\in V$ the function
$r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}.$
Proposition 2.4.
For $a\in V$ the function $r_{a}^{k}$ has the reproducing property
$\left[r_{a}^{k},g\right]=g(a)$
in $\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(V)$.
Proof.
Note that
$\left[r_{a}^{k},g\right]$ $=\left[P_{\operatorname{alt}}q_{a}^{k},g\right]=\left[q_{a}^{k},P_{\operatorname{alt}}g\right]$
$=\left[q_{a}^{k},g\right]=g(a)$
where we used the definition of $r_{a}^{k}$, $P_{\operatorname{alt}}^{*}=P_{\operatorname{alt}}$ and the reproducing property of $q_{a}^{k}$ sated in Proposition 2.2. ∎
Analogous properties hold in the space of symmetric polynomials $\operatorname{\mathcal{P}}_{\operatorname{sym}}^{k}(V)$ but we are not going to use them.
For $a\in V$ we have the translation operator
$T_{a}:\operatorname{\mathcal{P}}^{k}(V)\to\operatorname{\mathcal{P}}^{k}(V)\qquad\mbox{given by}\qquad T_{a}f(x)=f(x+a).$
This operator has an adjoint $T_{a}^{*}:\operatorname{\mathcal{P}}^{k}(V)\to\operatorname{\mathcal{P}}^{k}(V)$. To characterize the adjoint we use the notation $F^{k}$ to truncate polynomials, if we have a polynomial $f(x)=\sum_{\beta}x^{\beta}$ then
$F^{k}(f(x))=\sum_{\beta:|\beta|\leq k}x^{\beta}.$
We also denote with $M_{f}g=f.g$ the multiplication of polynomials. Define the operator
$F^{k}M_{q_{a}^{k}}:\operatorname{\mathcal{P}}^{k}(V)\to\operatorname{\mathcal{P}}^{k}(V)$
which consists of multiplying by $q_{a}^{k}$ and discarding in the result the terms of degree greater than $k$.
Proposition 2.5.
For $a\in V$ and $k\in\operatorname{\mathbb{N}}_{0}$ we have
$T_{a}^{*}=F^{k}M_{q_{a}^{k}}.$
Proof.
For $f,g\in\operatorname{\mathcal{P}}^{k}(V)$
$\left[F^{k}M_{q_{a}^{k}}f,g\right]$ $=e^{\langle\partial,a\rangle}f(\partial)g(x)|_{x=0}$
$=f(\partial)e^{\langle\partial,a\rangle}g(x)|_{x=0}$
$=f(\partial)g(x+a)|_{x=0}$
$=f(\partial)T_{a}g(x)|_{x=0}$
$=\left[f,T_{a}g\right].$
An alternative proof can be given by checking the equation in monomials $f(x)=x^{\beta}$ and $g(x)=x^{\beta}$ using the binomial formula for several variables. We have for multi-indices $\beta\leq\gamma$
$\left[x^{\beta},T_{a}x^{\gamma}\right]$ $=\left[x^{\beta},(x+a)^{\gamma}\right]=\left[x^{\beta},\sum_{\delta\leq\gamma}\binom{\gamma}{\delta}x^{\delta}a^{\beta-\delta}\right]$
$=\left[x^{\beta},\binom{\gamma}{\beta}x^{\beta}a^{\gamma-\beta}\right]=a^{\gamma-\beta}\beta!\frac{\gamma!}{\beta!(\gamma-\beta)!}.$
One the other hand, with a slight abuse of notation when using the not truncated exponential, we have
$\left[e^{\langle a,x\rangle}x^{\beta},x^{\gamma}\right]=\left[\sum_{\delta}\frac{1}{\delta!}a^{\delta}x^{\delta+\beta},x^{\gamma}\right]=\left[\frac{1}{(\gamma-\beta)!}a^{\gamma-\beta}x^{\gamma},x^{\gamma}\right]=a^{\gamma-\beta}\frac{\gamma!}{(\gamma-\beta)!}.$
The creation and annihilation operator are fundamental in the theory of Segal-Bargman space. They consist of multiplication by a variable and differentiation by the same variable. In the context of spaces of polynomials we define for $\alpha\in V$ the multiplication operator
$M_{\alpha}:f(x)\operatorname{\mathcal{P}}^{k}(V)\to\langle\alpha,x\rangle f(x)\operatorname{\mathcal{P}}^{k}(V),\qquad g(x)\mapsto\langle\alpha,x\rangle g(x)$
and the derivative operator
$A_{\alpha}:\langle\alpha,x\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)\to f(x)\operatorname{\mathcal{P}}^{k}(V),\qquad g(x)\mapsto\langle\alpha,\partial\rangle g(x).$
Proposition 2.6.
For $k\in\operatorname{\mathbb{N}}_{0}$ and $\alpha\in V$
$M_{\alpha}^{*}=A_{\alpha}.$
Proof.
For $g$ in the domain of $M_{\alpha}$ and $h$ in the domain of $A_{\alpha}$ we have
$\left[M_{\alpha}g,h\right]$ $=g(\partial)\langle\alpha,\partial\rangle h(x)|_{x=0}$
$=g(\partial)A_{\alpha}h(x)|_{x=0}$
$=\left[g,A_{\alpha}h\right],$
where we used the definition of the apolar inner product. ∎
The inverse of the annihilation and creation operators are the division operator $D_{\alpha}=M_{\alpha}^{-1}$ and the anti-derivative operator $I_{\alpha}=A_{\alpha}^{-1}$. They are adaptations of the adjoints of the divided difference or Bernstein-Gelfand-Gelfand operators introduced in Section $6$ of [PS09]. For $k\in\operatorname{\mathbb{N}}_{0}$, $\alpha\in V$ and a polynomial $f\in\operatorname{\mathcal{P}}(V)$ we define the division operator
$D_{\alpha}:\langle x,\alpha\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)\to f(x)\operatorname{\mathcal{P}}^{k}(V)$
as
$g(x)\mapsto D_{\alpha}g(x)=\frac{1}{\langle\alpha,x\rangle}g(x),$
and the anti-derivative operator
$I_{\alpha}:f(x)\operatorname{\mathcal{P}}^{k}(V)\to\langle x,\alpha\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)$
as
$g(x)\mapsto I_{\alpha}g(x)=\int^{\langle x,\alpha^{\circ}\rangle}_{0}g(x-\alpha t)dt,$
where $\alpha^{\circ}=\alpha/\langle\alpha,\alpha\rangle$. Note that the integration is done on the segment from $x$ to the projection of $x$ to the hyperplane $\{\alpha\}^{\perp}$. For $\alpha$ such that $\|\alpha\|=1$, the signed distance of $x$ to the hyperplane $\{\alpha\}^{\perp}$ is $\langle x,\alpha\rangle$. The proof of the following proposition is similar to the proof of Lemma $6.3$ in [PS09].
Proposition 2.7.
For the operators defined above we have
$D_{\alpha}^{*}=I_{\alpha},$
that is, the adjoint of the division operator is the anti-derivative.
Proof.
For a constant $c\neq 0$ and unit norm $\alpha\in V$ we have $D_{c\alpha}=\frac{1}{c}D_{\alpha}$. Also, by a linear change of variables in the definition of $I_{\alpha}$ we can assume that $\|\alpha\|=1$. Since an orthogonal change of coordinates leaves the equation unchanged we can prove the identity for the standard basis vector $\alpha=e_{1}$. We can also verify the identity for polynomials $x_{1}^{k}f(x_{2},\dots,x_{n})$ and $x_{1}^{l}g(x_{2},\dots,x_{n})$ since polynomials of this type span the spaces of polynomials. Since the operators $D_{e_{1}}$ and $I_{e_{1}}$ don’t act on the factors $f$ and $g$ and since
$\left[x_{1}^{k}f(x_{2},\dots,x_{n}),x_{1}^{l}g(x_{2},\dots,x_{n})\right]=\left[x_{1}^{k},x_{1}^{l}\right]\left[f(x_{2},\dots,x_{n}),g(x_{2},\dots,x_{n})\right]$
it is enough to verify the identity in the monomials $x_{1}^{k}$ and $x_{1}^{l}$. We have
$D_{e_{1}}(x_{1}^{k})=\frac{1}{x_{1}}x_{1}^{k}=x_{1}^{k-1}$
and
$I_{e_{1}}(x_{1}^{l})=\int_{0}^{x_{1}}t^{l}dt=\frac{1}{l+1}x_{1}^{l+1}.$
It is easy to very that $\left[x_{1}^{k-1},x_{1}^{l}\right]=\left[x_{1}^{k},\frac{1}{l+1}x_{1}^{l+1}\right]$. If $k\neq l+1$ then both side are zero and if $k=l+1$ then both sides are equal to $l!$. ∎
Until the end of this section we consider vector spaces which are the Cartan algebras $\operatorname{\mathfrak{t}}$ of Lie algebras of compact semi-simple Lie groups. One of the essential properties of the discriminant $\Delta:\operatorname{\mathfrak{t}}\to\operatorname{\mathbb{R}}$ is that it skew with respect to the action of $\operatorname{\mathcal{W}}$, $\Delta(w(x))=\operatorname{\epsilon}(w)\Delta(x)$ for $w\in\operatorname{\mathcal{W}}$. This follows from the fact that if $\alpha$ is a simple root the reflection through the plane $\langle\alpha,x\rangle=0$ sends $\alpha\mapsto-\alpha$ and permutes the other positive roots. The next proposition is well known.
Proposition 2.8.
Every polynomial $f\in\operatorname{\mathcal{P}}^{k}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$ is divisible by $\Delta$.
Proof.
For each reflection $w_{\alpha}\in\operatorname{\mathcal{W}}$ we have $f(w_{\alpha}(x))=-f(x)$ so $f$ vanishes on the hyperplane $\langle\alpha,x\rangle=0$ and the linear polynomial $\langle\alpha,x\rangle$ divides $f$. Since the Lie algebra is semi-simple its root system is reduced, that is, the only scalar multiple of a root $\alpha$ is $\alpha$. Therefore, the positive roots are relatively prime, since they only divide each other if they are scalar multiples. Since for each positive root the polynomial $\langle\alpha,x\rangle=0$ divides $f$ the discriminant $\Delta$ divides the polynomial $f$. ∎
This proposition implies that
$\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})=\Delta.\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}}).$
Therefore we can define the division by the discriminant
$D_{\Delta}:\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\qquad\mbox{given by}\qquad D_{\Delta}f(x)=\frac{f(x)}{\Delta(x)}.$
We can define $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ for an ordering $\alpha_{1},\alpha_{2},\dots,\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots. The order of the composition
$\Delta(x)\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{1}}}\prod_{\alpha\neq\alpha_{1}}\langle\alpha,x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{2}}}\prod_{\alpha\neq\alpha_{1},\alpha_{2}}\langle\alpha,x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{3}}}\ldots$
$\ldots\xrightarrow{D_{\alpha_{l-2}}}\langle\alpha_{l-1},x\rangle\langle\alpha_{l},x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{l-1}}}\langle\alpha_{l},x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{l}}}\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})$
does not change the final outcome. The operator
$I_{\Delta}:\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}$
is defined by $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ for an ordering $\alpha_{1},\alpha_{2},\dots\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots $\Phi^{+}$.
Proposition 2.9.
The operator $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ is independent of the ordering of positive roots and it satisfies
$I_{\Delta}=D_{\Delta}^{*}.$
Proof.
Since the order of the composition $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ is irrelevant and since $I_{\alpha}=D_{\alpha}^{*}$ for all $\alpha\in\operatorname{\mathfrak{t}}$ by Proposition 2.7 the conclusion of the statement of the theorem follows. ∎
We denote with $\mathbf{1}$ the constant polynomial $\mathbf{1}(x)=1$.
Proposition 2.10.
The iterated anti-derivative $I_{\Delta}$ satisfies
$I_{\Delta}\mathbf{1}=\frac{\Delta}{\left[\Delta,\Delta\right]}.$
Proof.
We have $D_{\Delta}(\operatorname{\mathbb{R}}\Delta)=\operatorname{\mathbb{R}}\mathbf{1}$ and $D_{\Delta}(\{\operatorname{\mathbb{R}}\Delta\}^{\perp})=\{\operatorname{\mathbb{R}}\mathbf{1}\}^{\perp}$. Therefore $I_{\Delta}\mathbf{1}=D^{*}_{\Delta}\mathbf{1}=d\Delta$ for a $d\in\operatorname{\mathbb{R}}$. We have
$1=\left[\mathbf{1},\mathbf{1}\right]=\left[D_{\Delta}\Delta,\mathbf{1}\right]=\left[\Delta,I_{\Delta}\mathbf{1}\right]=\left[\Delta,d\Delta\right]=d\left[\Delta,\Delta\right].$
Hence $d=\left[\Delta,\Delta\right]^{-1}$ and the conclusion of the proposition follows. ∎
3. The projection of an orbital measure to the Cartan algebra
For $a\in\operatorname{\mathfrak{t}}$ let $\nu_{a}$ be the orbital measure on $\operatorname{\mathcal{O}}_{a}$ and define the measure $\mu_{a}$ on the Cartan algebra as $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ where $\operatorname{pr}:\operatorname{\mathfrak{g}}\to\operatorname{\mathfrak{t}}$ is the orthogonal projection. The measure $\mu_{a}$ was characterized in [Z16, O13, F15], it can also be characterized as a Duistermaat-Heckman measure in the context of Hamiltonian torus actions, see Section $5$ in the book [GLS96] and the references therein. We start with an equation satisfied by the measure $\mu_{a}$.
Proposition 3.1.
For $a\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ the measure $\mu_{a}$ satisfies
(2) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
Proof.
Since $\langle\operatorname{Ad}_{g}x,y\rangle=\langle\operatorname{pr}(\operatorname{Ad}_{g}x),y\rangle$ for $y\in\operatorname{\mathfrak{t}}$, $x\in\operatorname{\mathfrak{g}}$ and $g\in G$ it is easy to very that the measure $\mu_{a}=\operatorname{pr}_{*}(\nu_{a})$ satisfies
$\int_{G}e^{\langle\operatorname{Ad}_{g}a,x\rangle}dg=\int_{G}e^{\langle\operatorname{pr}(\operatorname{Ad}_{g}a),x\rangle}dg=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
Therefore, the orbital integral (1) implies that the measure $\mu_{a}$ satisfies
$\operatorname{\mathcal{F}}_{a}(x)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(x)}\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}=\int_{\operatorname{\mathfrak{t}}}e^{\langle c,x\rangle}d\mu_{a}(c).$
Theorem 3.2.
For $a\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ the measure $\mu_{a}$ is characterized by
(3) $\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right),$
for every $f\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$.
Proof.
We write $\eta=\left[\Delta,\Delta\right]/\Delta(a)$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ in $x$ of equation (2) of Proposition 3.1 then we have
$\eta\frac{r_{a}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}(x)d\mu_{a}(c).$
If for a polynomial $f\in\operatorname{\mathcal{P}}^{k}(\operatorname{\mathfrak{t}})$ we take the inner product of both sides with $f_{\operatorname{sym}}=P_{\operatorname{sym}}f$ then in the LHS we get
$\left[\eta D_{\Delta}r_{a}^{k+l},f_{\operatorname{sym}}\right]$ $=\eta.\left[D_{\Delta}r_{a}^{k+l},f_{\operatorname{sym}}\right]$
$=\eta.\left[r_{a}^{k+l},I_{\Delta}f_{\operatorname{sym}}\right]$
$=\eta.\left[q_{a}^{k+l},I_{\Delta}f_{\operatorname{sym}}\right]$
$=\eta.\operatorname{ev}_{a}\left(I_{\Delta}f_{\operatorname{sym}}\right)$
where we used Propositions 2.9 in the second equality, Lemma 2.3 in the third and Proposition 2.4 in the final equality. On the other hand, in the LHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}q_{c}^{k}d\mu_{a}(c),f_{\operatorname{sym}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\left[q_{c}^{k},f_{\operatorname{sym}}\right]d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\operatorname{ev}_{c}(f_{\operatorname{sym}})d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{sym}}(c)d\mu_{a}(c)$
$=\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$
where we used Proposition 2.2 in the second equality and the symmetry of the measure $\mu_{a}$ in the final equality. Hence, the conclusion of the theorem holds. ∎
Example 3.3.
Let us check that the measure $\mu_{a}$ characterized in this way is a probability measure. If we take as polynomial $f=\mathbf{1}$ then by Theorem 3.2 and Proposition 2.10
$\int_{\operatorname{\mathfrak{t}}}1d\mu(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(I_{\Delta}\mathbf{1}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\operatorname{ev}_{a}\left(\frac{\Delta}{\left[\Delta,\Delta\right]}\right)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)}\frac{\Delta(a)}{\left[\Delta,\Delta\right]}=1.$
Example 3.4.
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$ and that $I_{\Delta}=I_{\alpha}$. To characterize the measure $\mu_{a}$ we compute
$\int_{\operatorname{\mathfrak{t}}}f(c)d\mu_{a}(c)$ $=\eta\operatorname{ev}_{a}\left(I_{\Delta}P_{\operatorname{sym}}f\right)$
$=\eta\operatorname{ev}_{a}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}P_{\operatorname{sym}}f(x-\alpha t)dt\right)$
$=\eta\int^{\left\langle a,\frac{1}{2}\alpha\right\rangle}_{0}P_{\operatorname{sym}}f(a-\alpha t)dt$
$=\frac{2}{a_{2}-a_{1}}\int^{\frac{1}{2}(a_{2}-a_{1})}_{0}\frac{1}{2}\left(f(a-\alpha t)+f(-a+\alpha t)\right)dt.$
It is easy to see that the density of the measure $\mu_{a}$ is given by the indicator function in the interval $\left[a,-a\right]$ of $\operatorname{\mathfrak{t}}$. Since adjoint orbits are spheres, that is, $\operatorname{\mathcal{O}}_{a}=\operatorname{Ad}_{\operatorname{SU}(2)}(a)\simeq S^{2}$ this is the fact known to Archimedes that the projection of the surface measure of a sphere to an axis is given by an indicator function.
4. The radial part of the convolution of orbital measures
Let $\nu_{a}$ and $\nu_{b}$ be the orbital measures on $\operatorname{\mathcal{O}}_{a}$ and $\operatorname{\mathcal{O}}_{b}$ for $a,b\in\operatorname{\mathfrak{t}}$. We characterize the radial part $\nu_{a,b}$ of the convolution of measures $\nu_{a}*\nu_{b}$. This was done in previous research in [F19, IZ80, CMZ19] using mainly Fourier analytic techniques. We start with an equation satisfied by the radial measure $\nu_{a,b}$.
Proposition 4.1.
For $a,b\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ and $\Delta(b)\neq 0$ the radial measure $\nu_{a,b}$ satisfies
(4) $\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\frac{1}{\Delta(x)}$ $\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(a),x\rangle}\right)\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(b),x\rangle}\right)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)e^{\langle w(c),x\rangle}\right)d\nu_{a,b}(c).$
Proof.
By Proposition 2.1 in [F19] the radial measure $\nu_{a,b}$ of the convolution $\nu_{a}*\nu_{b}$ is given by the Fourier-Laplace transform as
$\operatorname{\mathcal{F}}_{a}(z)\operatorname{\mathcal{F}}_{b}(z)=\int_{\operatorname{\mathfrak{t}}}\operatorname{\mathcal{F}}_{c}(z)d\nu_{a,b}(c).$
If we express the Fourier-Laplace transform terms $\operatorname{\mathcal{F}}_{a},\operatorname{\mathcal{F}}_{b}$ and $\operatorname{\mathcal{F}}_{c}$ using the orbital integral formula (1), and if we multiply both sides by $\Delta(x)$ and divide both sides by $\left[\Delta,\Delta\right]$ then we obtain the equation (4) in the statement of the proposition. ∎
Theorem 4.2.
For $a,b\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ and $\Delta(b)\neq 0$ the radial part $\nu_{a,b}$ of the convolution $\nu_{a}*\nu_{b}$ of orbital measures is characterized by
(5) $\int_{\operatorname{\mathfrak{t}}}\frac{f_{\operatorname{alt}}(c)}{\Delta(c)}d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right).$
for every alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{alt}(\operatorname{\mathfrak{t}})$.
Proof.
We set $\varphi={{\left[\Delta,\Delta\right]}}/{\Delta(a)\Delta(b)}$ and $l=|\Phi^{+}|$. If we take the part of degree not greater than $k\in\operatorname{\mathbb{N}}_{0}$ of equation (4) we get
$\varphi.F^{k}M_{r_{a}^{k}}\frac{r_{b}^{k+l}(x)}{\Delta(x)}=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c).$
For an alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k}(\operatorname{\mathfrak{t}})$ we take the inner product of both sides with $f_{\operatorname{alt}}$. On the LHS we have
$\left[\varphi.F^{k}M_{r_{a}^{k}}D_{\Delta}r_{b}^{k+l},f_{\operatorname{alt}}\right]$ $=\varphi.\left[\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)F^{k}M_{q_{w(a)}^{k}}D_{\Delta}r_{b}^{k+l},f_{\operatorname{alt}}\right]$
$=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\varphi.\left[\epsilon F^{k}M_{q_{w(a)}^{k}}D_{\Delta}r_{b}^{k+l},f_{\operatorname{alt}}\right]$
$=\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\varphi.\left[D_{\Delta}r_{b}^{k+l},\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right]$
$=\varphi.\left[D_{\Delta}r_{b}^{k+l},\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right]$
$=\varphi.\left[r_{b}^{k+l},I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right]$
$=\varphi.\left[q_{b}^{k+l},I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right]$
$=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}f_{\operatorname{alt}}\right)\right)$
where we used the definition $r_{a}^{k}=P_{\operatorname{alt}}q_{a}^{k}$ in the first equality, Proposition 2.5 in the third, Proposition 2.9 in the fifth, Lemma 2.3 in the sixth and Proposition 2.2 in the final equality. On the other hand, on the RHS we get
$\left[\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}r_{c}^{k}(x)d\nu_{a,b}(c),f_{\operatorname{alt}}\right]$ $=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\left[r_{c}^{k}(x),f_{\operatorname{alt}}\right]d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}\operatorname{ev}_{c}(f_{\operatorname{alt}})d\nu_{a,b}(c)$
$=\int_{\operatorname{\mathfrak{t}}}\frac{1}{\Delta(c)}f_{\operatorname{alt}}(c)d\nu_{a,b}(c),$
where we used Proposition 2.4 in the second equality. Combining the result of taking the inner product on both sides we obtain the equation in the statement of the theorem. Since all symmetric polynomials are of the form $f_{\operatorname{alt}}/\Delta$ for an alternating polynomial $f_{\operatorname{alt}}$ and since the measure $\nu_{a,b}$ is symmetric we see that (5) completely characterizes the measure $\nu_{a,b}$. ∎
We can get a slight variation of the characterization in Theorem 4.2 of the radial measure $\nu_{a,b}$.
Corollary 4.3.
The radial part $\nu_{a,b}$ of the convolution $\nu_{a}*\nu_{b}$ is characterized by
$\int_{\operatorname{\mathfrak{t}}}g(c)d\nu_{a,b}(c)=\frac{\left[\Delta,\Delta\right]}{\Delta(a)\Delta(b)}\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{|\operatorname{\mathcal{W}}|}\sum_{w\in\operatorname{\mathcal{W}}}\epsilon(w)T_{w(a)}(\Delta.P_{\operatorname{sym}}g)f\right)\right)$
for every polynomial $g\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$.
Proof.
An alternating polynomial $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$ can be written as $f_{\operatorname{alt}}=\Delta.P_{\operatorname{sym}}g$ for a polynomial $g\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$. Making this substitution in formula (5) and noting that
$\int_{\operatorname{\mathfrak{t}}}(P_{\operatorname{sym}}h)(c)d\nu_{a,b}(c)=\int_{\operatorname{\mathfrak{t}}}h(c)d\nu_{a,b}(c)$
for any polynomial $h\in\operatorname{\mathcal{P}}(\operatorname{\mathfrak{t}})$ the corollary follows. ∎
Example 4.4.
Consider the case in which $G=\operatorname{SU}(2)$, and $\alpha\in\operatorname{\mathfrak{t}}\subseteq\operatorname{\mathfrak{su}}(2)$ is given by $\alpha=e_{2}-e_{1}$ where $\{e_{1},e_{2}\}$ is the standard basis. Note that $\alpha^{\circ}=\frac{1}{2}\alpha$. For $f_{\operatorname{alt}}\in\operatorname{\mathcal{P}}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})$
$\int_{\operatorname{\mathfrak{t}}}f_{\operatorname{alt}}(c)\frac{1}{\Delta(c)}d\nu_{a,b}(c)$ $=\varphi.\operatorname{ev}_{b}\left(I_{\Delta}\left(\frac{1}{2}\left(T_{a}f_{\operatorname{alt}}-T_{-a}f_{\operatorname{alt}}\right)\right)\right)$
$=\varphi.\operatorname{ev}_{b}\left(\int^{\langle x,\alpha^{\circ}\rangle}_{0}\frac{1}{2}\left((f_{\operatorname{alt}}(x+a-\alpha t)-f_{\operatorname{alt}}(x-a-\alpha t)\right)dt\right)$
$=\varphi.\int^{\frac{1}{2}(b_{2}-b_{1})}_{0}\frac{1}{2}\left(f_{\operatorname{alt}}(b+a-\alpha t)-f_{\operatorname{alt}}(b-a-\alpha t)\right)dt$
where $\varphi=\frac{2}{(a_{2}-a_{1})(b_{2}-b_{1})}$. If we make the identification $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ given by $(x,-x)\to x$ then one can verify that the density function $\phi$ of the measure $\nu_{a,b}$ is given when $0<b<a$ by
$\phi(c)=\begin{cases}\frac{c}{4ab}&\text{if $a-b\leq c\leq a+b$}\\ \frac{-c}{4ab}&\text{if $-a-b\leq c\leq-a+b$}\\ 0&\text{otherwise.}\end{cases}$
Acknowledgements
I thank Pablo Zadunaisky for discussions and for suggesting the article [PS09].